jACK CLIMBED UP THE BEANSTALK AT A UNIFORM RATE. AT 2PM HE WAS 1/6 OF THE WAY UP AND AT 4PM HE WAS 3/4 UP . WHAT FRACTIONAL PART OF THE ENTIRE BEANSTALK HAD HE CLIMBED BY 3PM?

Answer by edjones(8007)

3/4 - 1/6
= 9/12 - 2/12
= 7/12 in 2 hrs.
7/12 * 1/2 = 7/24 in 1 hr
.
1/6 + 7/24
= 4/24 + 7/24
=11/24 OF THE ENTIRE BEANSTALK HAD HE CLIMBED BY 3PM.
.
Ed